OpenStudy (anonymous):

in the interval 0 less than or equal to x greater than 2pi the solutions of the equation sin^2x=sin x are?

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

first we solve for sin(x) and then we have to solve for x.

OpenStudy (anonymous):

so let us call \[sin(x)=z\] so as not to confuse the algebra with the trig

OpenStudy (anonymous):

\[z^2=z\] \[z^2-z=0\] \[z(z-1)=0\] \[z+0\] or \[z=1\]

OpenStudy (anonymous):

typo should be \[z=0\] or \[z=1\]

OpenStudy (anonymous):

now replace z by \[sin(x)\] and solve \[sin(x)=0\] to get \[x=0\] because \[sin(0)=0\] or \[sin(x)=1\] to get \[x=\frac{\pi}{2}\] because \[sin(\frac{\pi}{2})=1\]

OpenStudy (anonymous):

from the cheat sheet although if you are taking trig you should memorize these

OpenStudy (anonymous):

how did you get from z(z-1)= 0 to z=0

OpenStudy (anonymous):

nevermind i got it, and so coult another answer be pi then?